OpenStudy (anonymous):

Which factor can you cancel from the numerator and denominator to simplify 12 minus 6 times x all over 5 times x minus 10.? [Hint: (a−x) = − (x − a)] A:6 over 5. B:5 over 6. C:x + 2 D:x − 2

OpenStudy (campbell_st):

\[\frac{ 12 - 6x}{5x - 10} = \frac{-6( x - 2)}{5(x - 2)}\] see if you can identify the common factor

OpenStudy (anonymous):

I got it Thank you (: